What is the wavelength of the matter-wave associated with a proton moving at 309 m/s?

the wavelength of proton matter-wave: ______m

What is the wavelength of the matter-wave associated with a 181 kg astronaut (including her spacesuit) moving at the same speed?

the wavelength of astronaut matter-wave: ____m

What is the wavelength of the matter-wave associated with Earthmoving along its orbit around the Sun?

the wavelength of Earth matter-wave: _____m
Answer #1

l=h m(proton)= 1.67X10 27 kg 1 mv by d= 6.6481034 (1•67x10*27)(309) t = 1,29X10-9m =h (181)(309) = 1.187X10-38m 6.64*10-34 mv